OpenStudy (anonymous):

What is the answer to this limit: lim n->infinity 1/1-n?

OpenStudy (phi):

is this \[\lim_{n \rightarrow \infty}\frac{ 1 }{ 1-n }\]

OpenStudy (phi):

if n gets very large, this is close to \[ \frac{1}{-n} \] (adding 1 to the -n hardly matters) now what is \[ -\frac{1}{1000} , -\frac{1}{10000000},...\] and so on, approaching it's easier to see in decimal form: \[ -0.001 , -0.0000001, -0.0000000001 ...\]

OpenStudy (anonymous):

so it is close to 0?

OpenStudy (phi):

yes, the expression gets close (but never reaches) 0. However, we say the LIMIT equals 0